OpenStudy (anonymous):

what is f(-5) if f(x)=|2x-1|+10 a.21 b.19 c.-1 d.-21

zepdrix (zepdrix):

\[\Large f(\color{royalblue}{x}) \quad=\quad \left|2(\color{royalblue}{x})-1\right|+10\] \[\Large f(\color{royalblue}{-5}) \quad=\quad \left|2(\color{royalblue}{-5})-1\right|+10\] You want to simplify everything inside of the absolute brackets first. Once you get it down to a single value, the bars are telling you to `change it to a positive number`.

OpenStudy (anonymous):

yea I got 11

zepdrix (zepdrix):

11 inside the bars? :o

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\Large f(-5)=|-11|+10 \qquad\to\qquad f(-5)=11+10\] Yah that looks good! :)

OpenStudy (anonymous):

21 so a

zepdrix (zepdrix):

yay sirryan \c:/